Circular No. 5768

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 M. P. Rupen and R. A. Sramek, National Radio Astronomy Observa-

tory; S. D. Van Dyk and K. W. Weiler, Naval Research Laboratory; and N. Panagia, Space Telescope Science Institute, communicate: “Two separate observing sessions of SN 1993J, made on Apr. 13 at 6 cm (4.86 GHz) with the Very Large Array in ‘B’ configuration, yield flux density values of 0.33 ± 0.08 and 0.28 ± 0.05 mJy. These measurements confirm the possible detection of radio emission at 6 cm on that same day recently reported by Strom (IAUC 5762). Such a value is also consistent with the model results reported by Panagia, Van Dyk, and Weiler on IAUC 5762. From the model, SN 1993J is expected to strengthen rapidly in this radio band and reach a peak of ∼ 30 mJy by May 20. Radio monitoring with the VLA is continuing on a nearly daily basis at this and other wavelengths in order to follow the supernova’s rapidly changing behavior.”

S. Radford, R. Neri, S. Guilloteau, and D. Downes, Institut de Radio Astronomie Millimétrique, Grenoble, write: “We obtained observations of SN 1993J at α = 9^{h}51^{m}19^{s}.03, δ = +69^{o}15′26′′.5 (equinox 1950.0; ± 0′′.5) with the IRAM interferometer. At 87 GHz, the flux was 11 mJy on Apr. 9 and 11, 13.5 mJy on Apr. 12, and 12.5 mJy on Apr. 15. At 110 GHz, the flux was 10 mJy on Apr. 16 (all values ± 2 mJy). These millimeter fluxes are less than those measured at 15 and 22 GHz at Cambridge and with the VLA, indicating non-thermal emission with a spectral index of ∼ –0.5 on Apr. 9. Within the 3-mm band, the spectral index cannot be derived reliably because of the flux uncertainties. For a uniform expansion at 10 000 km/s and a distance of 3 Mpc, the angular diameter of the remnant is ∼ 0′′.00008 by Apr. 15. Hence the brightness temperature at 87 GHz is 3 × 10^{8} K. The decrease in flux between cm and mm wavelengths shows the foreground ionized plasma (presumably the remnant of the presupernova circumstellar envelope) is optically thin at mm wavelengths. The nonthermal radio spectrum is cut off exponentially at lower frequencies by the opaque foreground plasma. The lack of dramatic time variation in the 3-mm flux during Apr. 9-16 is further evidence that the foreground plasma has become transparent in the mm range.”

Corrigendum. On IAUC 5743, line –4, for indentification read identification

1993 April 20 Daniel W. E. Green